Kingbird Corp. purchased machinery for $381,150 on May 1, 2020. It is estimated that it will have a useful life of 10 years, salvage value of $18,150, production of 290,400 units, and working hours of 25,000. During 2021, Kingbird Corp. uses the machinery for 2,650 hours, and the machinery produces 30,900 units.

From the information given, compute the depreciation charge for 2021 under each of the following methods. (Round intermediate calculations to 2 decimal places, e.g. 5.25 and final answers to 0 decimal places, e.g. 45,892.)

(a)

Straight-line

$enter a dollar amount

Entry field with correct answer

(b)

Units-of-output

$enter a dollar amount

Entry field with correct answer

(c)

Working hours

$enter a dollar amount

Entry field with correct answer

(d)

Sum-of-the-years'-digits

$enter a dollar amount

Entry field with incorrect answer now contains modified data

(e)

Declining-balance (use 20% as the annual rate)

$enter a dollar amount

Entry field with incorrect answer
Answer #1

From the information given, compute the depreciation charge for 2021 under each of the following methods. (Round intermediate calculations to 2 decimal places, e.g. 5.25 and final answers to 0 decimal places, e.g. 45,892.)

(a)

Straight-line

(381150-18150/10) = 36300
(b)

Units-of-output

363000/290400*30900 = 38625
(c)

Working hours

363000/25000/2650 = 38478
(d)

Sum-of-the-years'-digits

363000*10/55*4/12+363000*9/55*8/12 = 61600
(e)

Declining-balance (use 20% as the annual rate)

363000*20%*4/12+363000*80%*20%*8/12 = 62920